OpenStudy (anonymous):

Seven times a number minus the number is -48. Find the number. the number is -8 the number is 8 the number is 6

OpenStudy (anonymous):

Similar to the last one you posted. How would you write this as an equation?

OpenStudy (anonymous):

Not quite. They don't mean to subtract the number and then multiply by seven, they mean multiply by seven, then subtract the number. So instead of 7(x-x)=-48, you're looking for 7x-x=-48.

OpenStudy (anonymous):

it is easy .just you can translate the problem into the equation and then you find the solution. let take the number N.so 7 times the number is 7N.and the equation will be:7N-N=-48 so N=-8.

OpenStudy (anonymous):

7times-8=-56 -56--8=-48

jhonyy9 (jhonyy9):

7*(-8) =-56 - -8 =-48